# Opara: Exploiting Operator Parallelism for Expediting DNN Inference on GPUs

### v1

@@ 3.3 Operator Launcher
Resource- and Interference-Aware Operator Launch Algorithm. Launching operators with heavy resource demands first to the GPU is likely to cause resource fragmentation, hindering the GPU executions of subsequent operators. Moreover, the GPU can thus be blocked due to the non-preemptive feature of kernel execution [11]. To mitigate such a problem and maximize the GPU utilization, we design a heuristic algorithm in Alg. 2 to greedily prioritize launching the operators with the least amount of GPU resource demands. Accordingly, it can maximize the parallel executions of multiple operators within a single model. To further mitigate the performance interference among operators [8], we aim to overlap the execution of computation-intensive operators and memory-intensive operators, as classified by our offline-collected operator table. Specifically, it first initializes and maintains a priority queue $\mathcal{Q}$ of operators in resource- and interference-aware operator launch order (line $1$). It then retrieves all operators to be launched with an indegree of 00 in a list $\mathcal{L}$, which alternates between a non-empty list of memory-intensive operators $\mathcal{L}_{mem}$ and a non-empty list of compute-intensive operators $\mathcal{L}_{comp}$ (lines $2$-$4$). Each time the operator requiring the least amount of GPU resources (e.g., shared memory, threads, registers) is chosen from $\mathcal{L}$ and then put into the queue $\mathcal{Q}$ (lines $5$-$6$). In particular, the potential GPU blocking issue faced by remaining large operators is noncritical in our scenario, as the operator list $\mathcal{L}$ is dynamic and can be compensated for the upcoming small operators to be launched. Finally, the operator list $\mathcal{L}_{mem}$ and $\mathcal{L}_{comp}$ are continuously updated by adding new operators with an indegree of 00 (lines $7$-$16$).
@@ 4 Implementation of Opara
We implement a prototype of Opara with around $1,000$ lines of Python codes, which have been integrated into PyTorch 2.0 as a plug-in module. The source codes are currently publicly available on GitHub (https://github.com/icloud-ecnu/Opara). Specifically, we employ torch.fx.Graph as the computation graph for DNN models in Opara. Its Intermediate Representation (IR) allows us to schedule DNN operators directly in the Python environment. In more detail, we leverage the torch.cuda.set_stream() API in PyTorch to launch operators on the CUDA Streams. In particular, as DNN operators are executed asynchronously on multiple streams for parallelized execution, appropriate operator synchronizations are required to guarantee that the parallelized computation conforms to the operator dependencies. We implement the operator synchronizations using the event.record() and stream.wait_event(event) APIs. Finally, we use torch.cuda.graph(g) to generate a CUDA Graph that can execute DNN operators in parallel based on the CUDA Streams. In summary, we build our prototype of Opara only using the high-level APIs of PyTorch in a lightweight manner, rather than modifying the computation graph construction module as in Nimble [10]. Accordingly, Opara is non-intrusive to the DL framework, and it can stably work as long as the framework APIs are not updated.
@@ 5.1 Experimental Setup
Hardware Configuration and Workloads. We conduct our experiments on an NVIDIA A100-PCIe-40GB GPU with Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz, and an NVIDIA GeForce RTX 2080 SUPER-8GB GPU with an Intel(R) Core(TM) i9-10920X CPU @ 3.50GHz. We implement Opara using PyTorch 2.0, CUDA 11.7, and cuDNN 8.5.0, as a plug-in module of PyTorch, as discussed in Sec. 4. Our experiments employ representative DNN models, including Inception-v3 [5], GoogLeNet [14], BERT [15], and T5 [17].
Baselines and Metrics. We compare DNN inference performance achieved by Opara with that achieved by the stock PyTorch (with CUDA Graph disabled), default sequential CUDA Graph, and Nimble [10]. Specifically, Nimble transforms the computation graph into a bipartite graph and then identifies its maximum matching to determine an appropriate stream for each operator. In particular, we focus on four key metrics including DNN inference latency, SM efficiency, and GPU memory consumption, as well as DNN inference throughput. All the experiment results are averaged over $1,000$ runs.
Figure 5: (a) Relative speedup, (b) SM efficiency and peak memory consumption of GPUs running representative DNN models with batch size set as $1$ achieved by PyTorch, CUDA Graph, Nimble, and Opara operator scheduling mechanisms.
@@ 5.2 Effectiveness of Opara
End-to-end Inference Latency. We first examine whether Opara can accelerate the DNN inference. As shown in Fig. 5, Opara consistently outperforms the other three baselines with four representative DNN models [^3] As BERT and T5 are GPU memory-intensive, we execute the two models on the A100, while the other two models are executed on the RTX 2080 SUPER.. Specifically, Opara can achieve 1.85×1.85$\times$1.85 × to 4.18×4.18$\times$4.18 × speedup compared to the stock PyTorch. This is because Opara utilizes CUDA Graph to eliminate the operator launch and function call overhead. Opara surpasses the default CUDA Graph by up to 1.68×1.68$\times$1.68 ×, simply because of the parallel execution of DNN operators in Opara. Furthermore, Opara outperforms Nimble by up to 1.29×1.29$\times$1.29 × because it judiciously alternates the scheduling of different types of operators with the lowest GPU resource consumption for each kernel launch time. Moreover, Opara initiates enough streams to increase parallelism (e.g., $28$ streams with Opara versus $4$ streams with Nimble for GoogLeNet), thereby maximizing the operator parallelism.
GPU Utilization and Memory. To unveil the performance gains of Opara, we proceed to look into the GPU utilization (i.e., SM efficiency of GPUs) and memory consumption during the execution of DNN models. As shown in Fig. 5, Opara exhibits a similar improvement in GPU utilization compared to the three baselines as in Fig. 5. Specifically, Opara significantly improves the GPU utilization compared to the stock PyTorch, because Opara mitigates the scheduling overhead of the stock PyTorch. When compared with the default CUDA Graph, Opara increases the GPU utilization of Inception-v3, GoogLeNet, BERT, and T5 by $36\%$, $58\%$, $20\%$, and $19\%$, respectively. Such performance gains come from the parallelized execution of operators. When compared to Nimble, Opara boosts the GPU utilization by 1.01×1.01$\times$1.01 × to 1.42×1.42$\times$1.42 × mainly due to the two facts: (1) maximizing stream allocations in Opara can increase operator parallelism opportunities, and (2) optimizing the operator launch order in Opara further minimizes the GPU idle time. Furthermore, the parallel execution of operators requires an increased amount of data to reside in GPU memory simultaneously, thereby leading to a higher peak GPU memory consumption of Opara than that of sequential executions (i.e., PyTorch and the default CUDA Graph).
@@
Timeline of Operator Executions. To understand the efficacy of Opara, we further illustrate the operator execution timeline by taking a segment of inference process of GoogLeNet as an example. In particular, we leverage NVIDIA Nsight System CLI [^4] Nsight System CLI: https://docs.nvidia.com/nsight-systems/UserGuide/index.html to track the operator executions. As depicted in Fig. 6, we observe that there are $4$ parallelizable operators numbered from $1-4$ during the segment of inference process. CUDA Graph executes these operators sequentially in a stream, and accordingly only operators $4$ and $1$ are executed during the time period. The other two operators (operators $2$ and $3$) are forced to queue up, which leads to a long execution time. Though Nimble can parallelize operators in the order of $1$, $2$, $3$, and $4$, it only schedules $2$ operators on two streams, causing a long GPU idle time. In contrast, Opara prioritizes operator $4$ and initiates more streams than Nimble, so that operators $4$, $1$, and $2$ can be executed at the same time and the operator parallelism is maximized. Accordingly, Opara can achieve the shortest inference latency by exploring operator parallelism compared with CUDA Graph and Nimble.
@@
Effectiveness of Opara on Transformer-Based Models. We conduct experiments with T5 and BERT model, and Opara outperforms Nimble by $9.3\%$ for the T5 model as shown in Fig. 5. This is because Opara optimizes the launch order of operators in T5 and schedules them into $6$ streams compared with $3$ streams in Nimble. Moreover, the operator diversity in T5 offers Opara overlap the compute-intensive Arange operators and the memory-intensive To and Ones operators, as shown in Fig. 7(a). In contrast, Opara achieves similar operator launch order and the same number of streams as Nimble as depicted in Fig. 7(b). This is because the parallelizable operators of BERT are always the Embedding operators or the Sgemm operators, which reduces the opportunity of operator overlapping and launch order optimization. Accordingly, Opara achieves marginal performance gains for BERT compared with Nimble, yet 1.08×1.08$\times$1.08 × to 4.06×4.06$\times$4.06 × speedup compared to the stock PyTorch and CUDA Graph as shown in Fig. 5.
@@
Figure 8: Inference throughput of Inception-v3 with PyTorch, CUDA Graph, Nimble, and Opara by varying the batch size from $1$ to $32$.
Figure 8: Inference throughput of Inception-v3 with PyTorch, CUDA Graph, Nimble, and Opara by varying the batch size from $1$ to $32$.
@@
Figure 9: Relative speedup of Inception-v3 with PyTorch, CUDA Graph, Nimble, and Opara by varying the batch size from $1$ to $8$.
@@
Throughput under Different Batch Sizes. We further examine the improvement of Opara in terms of DNN inference throughput across varying batch sizes. As depicted in Fig. 9, we observe that Opara consistently surpasses the other three baselines by varying the batch size from $1$ to $32$. Nevertheless, the performance gains of Opara gradually diminish as the batch size increases. As an example, Opara outperforms the default CUDA Graph by 1.41×1.41$\times$1.41 × and 1.09×1.09$\times$1.09 × when the batch size is $1$ and $32$, respectively. This is because the amount of GPU resources occupied by a single operator increases when dealing with larger batch sizes, which results in less amount of GPU resources available for the execution of parallelized operators. The results above also show that maximizing the operator parallelism can expedite DNN inference while improving the GPU utilization.
Effectiveness of Opara on GPUs with Sufficient Resources. To validate the efficacy of Opara on GPUs equipped with powerful computing resources, we repeat experiments by comparing performance gains with the three baselines on the A100. As shown in Fig. 9, we observe that Opara consistently outperforms the three baselines by varying the batch size from $1$ to $8$, mainly due to the better performance of operator parallelism on the A100. In more detail, both Opara and Nimble can benefit from the operator parallelism compared to the stock PyTorch and default CUDA Graph. Moreover, Opara can achieve 1.08×1.08$\times$1.08 × to 1.15×1.15$\times$1.15 × speedup compared to Nimble on the A100, which are slightly smaller performance gains compared to that on the RTX 2080 SUPER. This is because sufficient resources are available on the A100, resulting in GPU under-utilization of DNN inference and limited optimization opportunities for the operator launch order. As the batch size increases to $8$, a higher GPU resource demands of the operator enlarge the optimization space for adjusting the operator launch order, leading to higher performance gains of Opara compared to Nimble.
@@ 5.3 Runtime Overhead of Opara
We evaluate the runtime overhead of Opara in terms of algorithm computation time and inference profiling overhead. We first compare the computation time of the stream allocation algorithm in Opara and Nimble [10]. As listed in Table I, Opara incurs algorithm computation time of $0.58$ ms, $0.27$ ms, $0.5$ ms, and $2.8$ ms for BERT, GoogLeNet, Inception-v3, and T5, respectively. In contrast, such computation time in Nimble is $20.8$ ms, $5.80$ ms, $14.40$ ms, and $161.4$ ms for the four models, respectively, which is at least a magnitude larger than that of Opara. This is because Nimble requires a graph transformation together with an exhaustive search in the bipartite graph. Such a process is time-consuming with a time complexity in the order of $\mathcal{O}({n}^{3})$, where $n$ is the number of operators in a model DAG. In contrast, the time complexity of Opara can be reduced to the order of $\mathcal{O}({n})$. This is because the inner loop of the stream allocation algorithm in Opara only depends on the maximum width of the computation graph, which is typically small (i.e., below $20$). As DNN models become increasingly complex and memory-intensive in future [18], the number of operators will also grow exponentially. Such an algorithm overhead in Nimble becomes unacceptable when $n$ is large enough. In addition, as the Model Profiler needs to run the DNN inference only once, Opara requires several (i.e., $4.25$) milliseconds of profiling overhead in our experiment. In sum, the runtime overhead of Opara is practically acceptable.
TABLE I: Computation time (in milliseconds) of the stream allocation algorithm (i.e., Alg. 1) in Opara and Nimble with different DNN models.
BERT GoogLeNet Inception-v3 T5
Opara $0.58$ $0.27$ $0.50$ $2.8$
Nimble [10] $20.8$ $5.80$ $14.40$ $161.4$
@@ 6 Related Work
Inter-operator Parallelism within A Single Model. To parallelize the execution of DNN operators, Rammer [19] proposes the concept of “rtask” to allow fine-grained computing scheduling on a GPU device. It enables task combinations in different operators executed on SMs. To maximize the operator parallelism, Cocktailer [20] further proposes to co-scheduling control flow operators and data flow operators into a GPU device based on Rammer. Nimble [10] leverages the bipartite graph algorithm to adequately launch operators on CUDA streams. A recent work (i.e., IOS [12]) deploys operator fusion and dynamic programming to determine operator parallelization plans. However, Nimble and IOS require a lengthy search process and neglect the optimization space of operator launch order as well as the operator launch overhead. In contrast, Opara utilizes the CUDA Graph to eliminate such performance overhead. It also employs a lightweight stream allocation algorithm to achieve inter-operator parallelism. To reduce the GPU idle time and interference, Opara determines a feasible operator launch order according to operator resource demands.
[…]
Intra-operator Parallelism. Existing DL frameworks, such as PyTorch and TensorFlow, employ expert-optimized operator libraries to accelerate individual operators. To reduce human interactions, TVM [24] uses machine learning methods to automatically search for efficient codes. Such a search process is time-consuming and requires artificially-specified parameter space. To achieve fully automated code generation, Ansor [25] implement an automatic search space construction. As the number of compute units in commodity GPUs increases, an individual DNN operator cannot fully utilize the GPU resources. Accordingly, the stream allocation and operator launch algorithms in Opara can work with the intra-operator parallelism methods above to further improve the GPU resource utilization.

### v2

@@ 7             Update $indegree$ of ${s}\leftarrow$ $indegree$ of ${s}-1$;
Resource- and interference-aware operator launch algorithm. Launching operators with heavy resource demands first to the GPU is likely to cause resource fragmentation, hindering the GPU executions of subsequent operators. Moreover, the GPU can thus be blocked due to the non-preemptive feature of kernel execution [12]. To maximize the GPU utilization, the key idea of Alg. 2 is to greedily prioritize launching the operators with the least amount of GPU resource demands, aiming at maximizing the parallel executions of multiple operators within a model. To further mitigate the performance interference among operators [8], we simply overlap the execution of compute-intensive operators and memory-intensive operators, as classified by our offline-collected operator table.
Specifically, it first initializes and maintains a priority queue $\mathcal{Q}$ of operators in resource- and interference-aware operator launch order (line $1$). It then retrieves all operators to be launched with an indegree of 00 in a list $\mathcal{L}$, which alternates between the non-empty lists for memory-intensive operators $\mathcal{L}_{mem}$ and for compute-intensive operators $\mathcal{L}_{comp}$ (lines $2$-$4$). Each time the operator requiring the least amount of GPU resources (e.g., shared memory, threads, registers) is chosen from $\mathcal{L}$ and then put into the queue $\mathcal{Q}$ (lines $5$-$6$). In particular, the potential GPU blocking issue faced by the remaining large operators is noncritical in our scenario, as $\mathcal{L}$ is dynamic and can be compensated for the upcoming small operators to be launched. Finally, $\mathcal{L}_{mem}$ and $\mathcal{L}_{comp}$ are continuously updated by adding new operators with an indegree of 00 (lines $7$-$13$).
@@ 4 Implementation of Opara
We implement a prototype of Opara with around $1,000$ lines of Python codes, which have been integrated into PyTorch 2.0 as a plug-in module. The source codes are currently publicly available on GitHub (https://github.com/icloud-ecnu/Opara). Specifically, we employ torch.fx.Graph as the computation graph for DNN models in Opara. Its Intermediate Representation (IR) allows us to schedule DNN operators directly in Python. In more detail, we leverage the torch.cuda.set_stream() API in PyTorch to launch operators on the CUDA Streams. To particularly guarantee the operator dependency in parallelized executions of streams, we add the appropriate synchronization operators to the model graph using the event.record() and stream.wait_event(event) APIs. Finally, we use torch.cuda.graph(g) to generate a CUDA Graph that can execute DNN operators in parallel based on the CUDA Streams. In summary, we build our prototype of Opara only using the high-level APIs of PyTorch in a lightweight and non-intrusive manner, rather than modifying the computation graph construction module as in Nimble [11].
@@ 5.1 Experimental Setup
Hardware configuration and workloads. We conduct our experiments on an NVIDIA A100-PCIe-40GB GPU and an NVIDIA GeForce RTX 2080 SUPER-8GB GPU. We implement Opara based on CUDA 11.7, cuDNN 8.5.0, and as a plug-in module of PyTorch 2.0. Our experiments employ $6$ representative DNN models, including Inception-v3 [5], GoogLeNet, DeepFM [18], NASNet [^5] https://huggingface.co/timm/nasnetalarge.tf_in1k, BERT, and T5666https://huggingface.co/google-t5. The first three models are executed on the RTX 2080 GPU, while the latter three are run on the A100 GPU.
Baselines and metrics. We compare DNN inference performance of Opara with that of the stock PyTorch (with CUDA Graph disabled), default sequential CUDA Graph, ONNX Runtime (with operator fusion enabled), Rammer [9], and Nimble [11]. Specifically, we implement Rammer’s BFS-based operator scheduling algorithm (named Wavefront) into PyTorch 2.0. Nimble transforms the computation graph into a bipartite graph and identifies its maximum matching to determine an appropriate stream for each operator. In particular, we focus on $4$ key metrics including DNN inference latency, SM efficiency, and GPU memory consumption, as well as DNN inference throughput. All the experiment results are averaged over $1,000$ runs.
Figure 5: (a) Relative speedup, (b) SM efficiency and peak memory consumption of GPUs running representative DNN models with batch size set as $1$ achieved by PyTorch, CUDA Graph, ONNX Runtime, Rammer, Nimble, and Opara operator scheduling mechanisms.
@@ 5.2 Effectiveness of Opara
End-to-end inference latency. As shown in Fig. 5, Opara consistently outperforms the five baselines for six representative DNN models. Specifically, Opara can achieve 1.80×1.80$\times$1.80 × to 10.97×10.97$\times$10.97 × speedup compared to the stock PyTorch. This is because Opara utilizes CUDA Graph to eliminate the operator launch and function call overhead. Opara surpasses the default CUDA Graph by up to 1.68×1.68$\times$1.68 ×, simply because of the parallel execution of DNN operators in Opara. Though the operator fusion technique outperforms the stock PyTorch, Opara achieves a higher speedup by up to 2.97×2.97$\times$2.97 × and 1.18×1.18$\times$1.18 ×, compared with ONNX Runtime and Rammer, respectively. Such performance improvements above are mainly due to two facts: First, the operator fusion cannot combine all parallelizable operators based on the pre-defined fusion rules, which cannot fully utilize the GPU resources. Second, Opara accelerates model inference through operator parallelization, while the Wavefront scheduling algorithm in Rammer introduces additional synchronization overhead (i.e., the unnecessary operator waiting time during wave executions). Furthermore, Opara outperforms Nimble by up to 1.29×1.29$\times$1.29 × because it judiciously alternates the scheduling of different types of operators with the lowest GPU resource consumption for each kernel launch time. Moreover, Opara initiates enough streams to increase parallelism (e.g., $28$ streams with Opara versus $4$ streams with Nimble for GoogLeNet), thereby maximizing the operator parallelism.
GPU utilization and memory consumption. To unveil the performance gains of Opara, we proceed to look into the GPU utilization (i.e., SM efficiency) and memory consumption during the model inference. As shown in Fig. 5, Opara exhibits a similar improvement in GPU utilization compared to the five baselines as in Fig. 5. Specifically, Opara significantly improves the GPU utilization compared to the stock PyTorch, because Opara mitigates the scheduling overhead of the stock PyTorch. When compared with the default CUDA Graph, Opara increases the GPU utilization of Inception-v3, GoogLeNet, DeepFM, NASNet, BERT, and T5 by $36\%$, $58\%$, $126\%$, $48\%$, $20\%$, and $19\%$, respectively. Such performance gains mainly come from the parallelized execution of operators. When compared to ONNX Runtime, Rammer, and Nimble, Opara boosts the GPU utilization by up to 3.86×3.86$\times$3.86 ×, 1.36×1.36$\times$1.36 × and 1.42×1.42$\times$1.42 × mainly because (1) maximizing stream allocations in Opara can increase operator parallelism opportunities, and (2) optimizing the operator launch order in Opara further minimizes the GPU idle time. Furthermore, the parallel execution of operators requires an increased amount of data to reside in the GPU memory simultaneously, thereby leading to a higher peak GPU memory consumption of Opara than that of sequential executions.
@@
Timeline of operator executions. We further illustrate the operator execution timeline by taking a segment of inference process of GoogLeNet as an example. In particular, we leverage NVIDIA Nsight System CLI [^7] https://docs.nvidia.com/nsight-systems/UserGuide/index.html to track the timeline of operator executions. As depicted in Fig. 6, we observe that CUDA Graph executes the $4$ operators sequentially in a stream, and only operators $4$ and $1$ appear within the time window. The remaining operators $2$ and $3$ are forced to queue up, which leads to a long inference time. Though Nimble can parallelize operators in the order of $1$, $2$, $3$, and $4$, it only schedules $2$ operators on two streams, causing a long GPU idle time. In contrast, Opara prioritizes operator $4$ and initiates more streams than Nimble, so that operators $4$, $1$, and $2$ can be executed in parallel to maximize the operator parallelism. Accordingly, Opara can achieve the shortest inference latency by exploring operator parallelism compared with CUDA Graph and Nimble.
@@
Effectiveness of Opara on Transformer-based models. We conduct experiments with T5 and BERT model, and Opara outperforms Nimble by $9.3\%$ for the T5 model as shown in Fig. 5. This is because Opara optimizes the launch order of operators in T5 and schedules them into $6$ streams compared with $3$ streams in Nimble. Moreover, the operator diversity in T5 offers Opara overlap the compute-intensive Arange operators and the memory-intensive To and Ones operators, as shown in Fig. 7(a). For BERT, however, Opara achieves a similar operator launch order and the same number of streams as Nimble as depicted in Fig. 7(b). This is because the parallelizable operators of BERT are always the Embedding operators or the Sgemm operators, which reduces the opportunity for operator overlapping and launch order optimization. Accordingly, Opara achieves marginal performance gains for BERT compared with Nimble, yet 1.08×1.08$\times$1.08 × to 4.06×4.06$\times$4.06 × speedup compared to the stock PyTorch and CUDA Graph as shown in Fig. 5.
@@
Figure 8: Inference throughput of Inception-v3 with Opara and the five baselines by varying the batch size from $1$ to $32$ on an RTX 2080 GPU.
@@
Figure 9: Relative speedup of Inception-v3 with Opara and the five baselines by varying the batch size from $2$ to $32$ on an A100 GPU.
@@
Throughput under different batch sizes. As depicted in Fig. 9, we observe that Opara consistently surpasses the five baselines except for ONNX Runtime by varying the batch size from $1$ to $32$. Nevertheless, the performance gains of Opara gradually diminish as the batch size increases. As an example, Opara outperforms the default CUDA Graph by 1.41×1.41$\times$1.41 × and 1.09×1.09$\times$1.09 × when the batch size is $1$ and $32$, respectively. This is because the amount of GPU resources occupied by a single operator increases when dealing with larger batch sizes, resulting in fewer GPU resources available for the execution of parallelized operators. This also explains why Opara exhibits marginal throughput improvement for large batch sizes of $16$ and $32$. The results above also show that maximizing the operator parallelism can also improve the inference throughput.
Effectiveness of Opara on high-end GPUs with sufficient resources. We repeat the inference experiment of Inception-v3 on a high-end GPU (i.e., A100). As shown in Fig. 9, we observe that Opara consistently outperforms the five baselines by varying the batch size from $2$ to $32$, mainly because operator parallelism works well for high-end GPUs with sufficient resources. In more detail, Opara achieves an inference speedup by up to 2.08×2.08$\times$2.08 ×, 1.29×1.29$\times$1.29 ×, and 1.15×1.15$\times$1.15 × compared to ONNX Runtime, Rammer, and Nimble, respectively. In particular, Opara achieves speedups of 1.47×1.47$\times$1.47 × and 1.18×1.18$\times$1.18 × relative to ONNX Runtime for batch sizes of $16$ and $32$, which is larger than the results achieved on the RTX 2080 GPU. This is because the A100 GPU provides sufficient resources, which allows the operator parallelism to achieve more performance gains than the operator fusion.
@@ 5.3 Runtime Overhead of Opara
We evaluate the runtime overhead of Opara in terms of algorithm computation time and inference profiling overhead. As listed in Table I, Opara can reduce the computation time of the stream allocation algorithm by up to two orders of magnitude compared with Nimble [11]. This is because Nimble requires a graph transformation together with an exhaustive search in the bipartite graph. Such a process is time-consuming with a complexity in the order of $\mathcal{O}({n}^{3})$, where $n$ is the number of operators in a model DAG. In contrast, the time complexity of Opara can be reduced to the order of $\mathcal{O}({n})$, simply because the inner loop of Alg. 1 (lines $3$-$10$) in Opara only depends on the maximum width (i.e., typically below $20$) of the computation graph. Accordingly, as DNN models become increasingly complex [19], the number of operators $n$ gets even larger, while the algorithm computation overhead of Opara can still be well contained. In addition, as the Model Profiler needs to run the DNN inference only once, Opara requires several (i.e., $4.25$) milliseconds of profiling overhead in our experiment. In sum, the runtime overhead of Opara is practically acceptable.
TABLE I: Computation time (in milliseconds) of the stream allocation algorithm in Opara (i.e., Alg. 1) and Nimble [11] for various models.
BERT GoogLeNet NASNet Inception-v3 T5
Opara $0.58$ $0.27$ $1.75$ $0.50$ $2.8$
Nimble $20.8$ $5.80$ $257.83$ $14.40$ $161.4$
@@ 6 Related Work
Inter-operator parallelism within a single model. To parallelize the execution of DNN operators, Rammer [9] proposes fine-grained operator scheduling based on the Wavefront algorithm and enables operator fusion on a GPU device. To increase the operator parallelism, Cocktailer [20] further co-schedules control flow and data flow operators based on Rammer. The two prior works above operate at the compilation level, which requires significant compilation overhead and manual customization of operators. Orthogonal to them, Opara focuses on the runtime operator scheduling optimization of the stream allocation and the operator launch order. A recent work Nimble [11] leverages the bipartite graph algorithm to schedule operators on CUDA streams adequately. IOS [13] deploys operator fusion and dynamic programming to determine operator parallelization plans. However, Nimble and IOS require a lengthy search process and neglect the optimization space of operator launch order. In contrast, Opara utilizes the CUDA Graph to eliminate such performance overhead. It also employs a lightweight stream allocation algorithm to achieve inter-operator parallelism. To reduce the GPU idle time and interference, Opara determines a feasible operator launch order according to operator resource demands.
Inter-operator parallelism among different models. To improve GPU utilization, several works parallelize operators from multiple models co-located on a GPU device. For example, S3DNN [21] and Abacus [22] optimize the co-location of operators from different models and schedule them to the corresponding steams. To minimize the model co-location interference, iGniter [8] and Orion [15] focus on optimizing the GPU resource allocation and operator scheduling on multiple prioritized streams, respectively. Paella [16] dispatches the optimal kernel from multiple models by jointly considering the remaining time and model fairness. Different from optimizing the inference co-location, Opara minimizes the inference latency while increasing the GPU utilization by parallelizing operators within a single model. Moreover, it achieves inter-operator parallelism as a plug-in module of PyTorch 2.0 without developing a new DL runtime or framework.
Intra-operator parallelism. Existing DL frameworks, such as PyTorch and TensorFlow, employ expert-optimized operator libraries to accelerate individual operators. TVM [23] uses machine learning methods to automatically search for efficient operators, which is time-consuming and requires the specified parameter space manually. To achieve automated code generation, Ansor [24] implements an automatic search space construction. As a single DNN operator cannot fully utilize GPU resources in general, Opara can work with the intra-operator parallelism methods above to further improve GPU resource utilization.